# Chow-Lefschetz motives

### v6

@@ Introduction
In two fundamental papers [25, 26], Milne put together various “Lefschetz groups” which had been attached to abelian varieties in the literature, showing that they assemble to form a Tannakian group associated to a category $\operatorname{\mathbf{LMot}}(k)$ of “Lefschetz motives” over an algebraically closed field $k$ [^1] To avoid confusions with the Lefschetz motive and Lefschetz decompositions, and also to honour Milne’s invention, it might be better to change the name “Lefschetz motives” to “Milne motives”. We shall refrain from it here, but hope this suggestion will be taken up by other mathematicians. This rigid $\otimes$-category has remarkable properties: homological and numerical equivalences agree for any Weil cohomology, and the fixed points of the Lefschetz group of $A$ on the cohomology of an abelian $k$-variety $A$ consist of the subring generated by divisor classes. One can specialise Lefschetz motives of CM abelian varieties from characteristic 00 to characteristic $p$, which allowed Milne to prove his famous theorem: the Hodge conjecture for CM abelian varieties over $\mathbf{C}$ implies the Tate conjecture for abelian varieties over finite fields [25, Th. 7.1]. André used Milne’s method in [3] to prove an unconditional result: all Tate cycles over an abelian variety in positive characteristic are “motivated” in his sense. Three questions arise from Milne’s construction:
@@
    Does $\operatorname{\mathbf{LMot}}(k)$ make sense for other adequate equivalence relations than homological and numerical equivalences? This question is raised in [26, p. 671].
@@
    Can one define $\operatorname{\mathbf{LMot}}(k)$ when $k$ is not algebraically closed? [^2] In [25, Introduction], it is announced that $\operatorname{\mathbf{LMot}}(k)$ is constructed over any field $k$, but [25, 1.4] refers to [26, 5.5] which assumes $k$ algebraically closed (the $k$ of [25] is the $\Omega$ of [26]).
@@
    Can one give a direct proof of Milne’s computation of the Tannakian group of his category in terms of Lefschetz groups, which follows in [26, §3] from a case-by-case analysis?
@@
The first two questions amount to asking whether sums of intersection products of divisor classes on abelian varieties are preserved under direct images by projections of the form $A\times B\to B$, where $A,B$ are abelian varieties. For $k$ algebraically closed and for homological equivalence, Milne proves this in [26, Cor. 5.5] as a consequence of his computation of Lefschetz groups. Here, we prove it in general. Namely, for an abelian variety $A$ over a field $k$, let $CH^{*}(A)$ denote the Chow ring of $A$, with rational coefficients, and let $L^{*}(A)$ be the subring of $CH^{*}(A)$ generated by $CH^{1}(A)$.
@@
As Beauville pointed out, the version of this theorem for numerical equivalence was proven previously by Schoen [32, Prop. 1.4], using a different method. See also Polishchuk [30, Th. 2.1].
@@
This theorem is in stark contrast with the case of usual motives, where algebraic and homological equivalences do not agree, by a famous theorem of Griffiths [14]. A more recent and relevant example, due to Schoen, is the cube of the Fermat elliptic curve $E$ in characteristic 00 [33, (0.2) and (14.1)]. As a consequence, $\dim_{\mathbf{Q}}\operatorname{Coker}(L^{*}(E^{3})\to CH^{*}(E^{3}))=+\infty$; this is remarkable, since this map becomes an isomorphism modulo numerical equivalence by [15] (see also [20]).
@@
Theorem 5 ([25, Prop. 1.8] when $k$ is algebraically closed).
@@
where $S$ is the set of isogeny classes of simple abelian $k_{s}$-varieties, and $\nu$ is induced by the $\nu_{A}$’s. (The corresponding groups $\operatorname{S\mathbb{U}}(A)$ were computed in [26, Table 2 p. 655].)
@@
Let $A$ be an elliptic curve. If $A$ has no complex multiplication, then $E=k$ and $\mathbb{U}(A)=\mathbf{GL}_{2}$. If $A$ has complex multiplication by $F$ in characteristic 00, then $\mathbb{U}(A_{E})=R_{F/\mathbf{Q}}\mathbb{G}_{m}$ and $E=kF$. If $A$ is ordinary in charactertistic $p$, then $E=k$, while if it is supersingular, then $E=kF_{0}$ if $k_{0}$ is the field of constants of $E$ and $F_{0}$ is the smallest (at most quadratic) extension of $k_{0}$ such that $\operatorname{End}^{0}(A_{E_{0}})$ is a quaternion algebra.
@@ 2. Technical lemmas
Since $CH^{1}_{(0)}(A)=L^{1}_{(0)}(A)$ generates $L^{*}_{(0)}(A)$ multiplicatively and since $CH^{1}_{(0)}(A)\to A^{1}_{\operatorname{num}}(A)$ is surjective, $\rho$ is surjective. On the other hand, O’Sullivan [29, Th. 6.1.1] has constructed a ring-theoretic section of the homomorphism $CH^{*}(A)\to A^{*}_{\operatorname{num}}(A)$ which sends $A^{1}_{\operatorname{num}}(A)$ into $CH^{1}_{(0)}(A)$, hence restricts to a ring-theoretic section $\sigma$ of $\rho$. But $CH^{1}_{(0)}(A)\to A^{1}_{\operatorname{num}}(A)$ is even an isomorphism, hence $\sigma$ is surjective. ∎
@@ 4.1. A crude category
When $f$ is a homomorphism, this follows from Theorem 1 since $[\gamma_{f}]=(\gamma_{f})_{*}([A])$. (See [25, 1.3] or [26, Th. 5.10] for a different proof.) In general, $f$ is the composition of a homomorphism and a translation, so it remains to handle the latter case. But if $a\in A(k)$, the translation defined by $a$ may be written as a composition
@@
As usual, we define $\operatorname{\mathbf{LMot}}_{\sim}^{\operatorname{eff}}(k)_{0}$ as the Karoubian envelope of $\operatorname{\mathbf{LCorr}}_{\sim}(k)_{0}$. As observed in [26, p. 672], the Chow-Künneth projectors of $A$ belong to $L^{*}(A\times A)$ as well as the Lefschetz isomorphisms of [24] and their inverses, which justifies the second claim already in $\operatorname{\mathbf{LMot}}_{\sim}^{\operatorname{eff}}(k)_{0}$. Note in particular that $\mathbb{L}\in\operatorname{\mathbf{LMot}}^{\operatorname{eff}}_{\sim}(k)_{0}$ by using an elliptic curve (which makes it unnecessary to involve projective spaces in its definition as in [25]), so that we can define $\operatorname{\mathbf{LMot}}_{\sim}(k)_{0}$ by $\otimes$-inverting $\mathbb{L}$. Its rigidity is checked on additive generators $Lh_{\sim}(A)$ in the usual way, using the diagonal of $A\times A$ to define unit and counit, and (4.2) is also checked as usual. ∎
@@ 4.4. Proof of Theorem 4
The existence of the Deninger-Murre Chow-Künneth projectors (Corollary 4.1) is the analogue of [28, I, Conj. (A)]. Reasoning as in [28, I, 2.5.4], the corresponding analogue to Murre’s filtration then verifies $F^{\nu}L^{j}(A)=\bigoplus_{s\geq\nu}L^{j}_{s}(A)$ for $0\leq\nu\leq j$ and $F^{j+1}L^{j}(A)=0$. Since $\operatorname{Ker}(L^{g}(A\times A)\to L^{g}_{\operatorname{num}}(A\times A))$ is a nilideal by [23], any other choice of Chow-Künneth projectors is conjugate to the canonical one under a self-correspondence $1+n$ with $n\sim_{\operatorname{num}}0$, as in [16, Lemma 5.4]. By (iii) and (iv) applied to $A\times A$, we have $n\in L^{g}_{(>0)}(A\times A)$, hence $nL^{j}_{(s)}(A)\subseteq\bigoplus_{t>s}L^{j}_{(t)}(A)$. This implies that the filtration $F^{\nu}L^{*}(A)$ does not depend on the choice of a Chow-Künneth decomposition, which is the analogue of Murre’s Conjecture (C). Finally, we get the analogue of Murre’s conjecture (B) (resp. (D)) from (iv) (resp. (iii)).
@@ 5.1. Centralisers in Tannakian categories
Recall [31, III.1] (see also [4, 5.3.1]) that one can define a $K$-linear Tannakian category $\mathcal{A}_{K}$ and a $Q$-linear $\otimes$-functor $(-)_{K}:\mathcal{A}\to\mathcal{A}_{K}$ such that $\mathcal{A}(A,B)\otimes_{Q}K\to\mathcal{A}_{K}(A_{K},B_{K})$ is an isomorphism for any $A,B\in\mathcal{A}$ and that $\omega$ extends canonically to a fibre functor on $\mathcal{A}_{K}$. This reduces us to the case where $Q=K$. The claim then follows from Tannakian theory, since $\mathcal{A}$ gets identified with $\operatorname{\mathbf{Rep}}_{K}(H)$. ∎
@@
$\rho^{2}(f)=\omega(u)^{-1}(1_{\omega(L)}\otimes\rho(f)^{\vee})\omega(u)\\ =\omega(u)^{-1}(1_{\omega(L)}\otimes(\omega(u)^{-1}(1_{\omega(L)}\otimes f^{\vee})\omega(u))^{\vee})\omega(u)\\ =\omega(u)^{-1}(1_{\omega(L)}\otimes(\omega(u)^{\vee}(1_{\omega(L^{\vee})}\otimes f)(\omega(u)^{-1})^{\vee})\omega(u)\\ =\omega(u^{-1}u^{\vee})f\omega((u^{-1})^{\vee}u)$
@@
where the fibre product is with respect to the second projections (cf. [26, Def. 4.6]). Here, we use the isomorphisms (5.2) for $rM$ and $M\oplus N$ obtained by direct sums from those of $M$ and $N$.
@@ 5.2. Weight gradings
is surjective for all $r,n$ by [26, Prop. 3.4] (whose proof uses invariant theory). The claim now follows from Lemma 5.1.
@@ 5.3. An “easy” exactness criterion
(In other words, Condition H0 is sufficient in [19, Lemma C.1] when $G$ and $H$ are proreductive.)
@@ 5.4. Action of a profinite group
See also [21, §5] for a different presentation.
@@
is exact at $G$ in general. This is true at least if $\mathcal{A}$ is semi-simple:
@@
In c), $\rho$ exists by [18, Prop. 5.3], which also gives the following recipe: for $S$ simple, $\rho(S)=S$ if $S\in\mathcal{A}^{0}$ and $\rho(S)=0$ otherwise. It suffices to check that this matches with the formula of the statement, which follows from b).
@@ 5.5. Proof of Theorem 5
Proposition 5.17 also applies to prove [10, Prop. 6.23 (a)] – except for the connectedness of $G(k_{s})$ which would require the “Hodge = absolute Hodge” conjecture –, and [2, 4.6, exemples]. It also applies to numerical Grothendieck motives under the standard conjecture D or, unconditionally, to this category restricted to numerical motives of abelian type in characteristic 00. Since the Tannakian group $G_{s}$ is not connected (see [26, Table 2 p. 655]), one sees that the expected connectedness of the motivic Galois group over separably closed fields cannot be hoped to be proven by purely formal arguments in the above style.
@@ 6. Remarks and questions
In particular, do the refined (Chow-)Künneth decompositions of [22, Th. 7.7.3]
@@
coincide in the two categories? This is true for $i=2$ by a generalisation of [22, Proof of Prop. 7.2.3]. However the answer is no in general, as pointed out by Peter O’Sullivan by the following simple argument:

### v7

@@ Introduction
In two fundamental papers [23, 24], Milne put together various “Lefschetz groups” which had been attached to abelian varieties in the literature, showing that they assemble to form a Tannakian group associated to a category $\operatorname{\mathbf{LMot}}(k)$ of “Lefschetz motives” over an algebraically closed field $k$ [^1] To avoid confusions with the Lefschetz motive and Lefschetz decompositions, and also to honour Milne’s invention, it might be better to change the name “Lefschetz motives” to “Milne motives”. We shall refrain from it here, but hope this suggestion will be taken up by other mathematicians. This rigid $\otimes$-category has remarkable properties: homological and numerical equivalences agree for any Weil cohomology, and the fixed points of the Lefschetz group on the cohomology of an abelian $k$-variety $A$ consist of the subring generated by divisor classes. One can specialise Lefschetz motives of CM abelian varieties from characteristic 00 to characteristic $p$, which allowed Milne to prove his famous theorem: the Hodge conjecture for CM abelian varieties over $\mathbf{C}$ implies the Tate conjecture for abelian varieties over finite fields [23, Th. 7.1]. André used Milne’s method in [3] to prove an unconditional result: all Tate cycles over an abelian variety in positive characteristic are “motivated” in his sense. Three questions arise from Milne’s construction:
@@
    Does $\operatorname{\mathbf{LMot}}(k)$ make sense for other adequate equivalence relations than homological and numerical equivalences? This question is raised in [24, p. 671].
@@
    Can one define $\operatorname{\mathbf{LMot}}(k)$ when $k$ is not algebraically closed? [^2] In [23, Introduction], it is announced that $\operatorname{\mathbf{LMot}}(k)$ is constructed over any field $k$, but [23, 1.4] refers to [24, 5.5] which assumes $k$ algebraically closed (the $k$ of [23] is the $\Omega$ of [24]).
@@
    Can one give a direct proof of Milne’s computation of the Tannakian group of his category in terms of Lefschetz groups, which follows in [24, §3] from a case-by-case analysis?
@@
The first two questions amount to asking whether sums of intersection products of divisor classes on abelian varieties are preserved under direct images by projections of the form $A\times B\to B$, where $A,B$ are abelian varieties. For $k$ algebraically closed and for homological equivalence, Milne proves this in [24, Cor. 5.5] as a consequence of his computation of Lefschetz groups. Here, we prove it in general. Namely, for an abelian variety $A$ over a field $k$, let $CH^{*}(A)$ denote the Chow ring of $A$, with rational coefficients, and let $L^{*}(A)$ be the subring of $CH^{*}(A)$ generated by $CH^{1}(A)$.
@@
As Beauville pointed out, the version of this theorem for numerical equivalence was proven previously by Schoen [29, Prop. 1.4], using a different method. See also Polishchuk [27, Th. 2.1].
@@
This theorem is in stark contrast with the case of usual motives, where algebraic and homological equivalences do not agree, by a famous theorem of Griffiths [13]. A more recent and relevant example, due to Schoen, is the cube of the Fermat elliptic curve $E$ in characteristic 00 [30, (0.2) and (14.1)]. As a consequence, $\dim_{\mathbf{Q}}\operatorname{Coker}(L^{*}(E^{3})\to CH^{*}(E^{3}))=+\infty$; this is remarkable, since this map becomes an isomorphism modulo numerical equivalence by [14] (see also [18]).
@@
Theorem 5 ([23, Prop. 1.8] when $k$ is algebraically closed).
@@
where $S$ is the set of isogeny classes of simple abelian $k_{s}$-varieties, and $\nu$ is induced by the $\nu_{A}$’s. (The corresponding groups $\operatorname{S\mathbb{U}}(A)$ were computed in [24, Table 2 p. 655].)
@@
Let $A$ be an elliptic curve. If $A$ has no complex multiplication, then $E=k$ and $\mathbb{U}(A)=\mathbf{GL}_{2}$. If $A$ has complex multiplication by $F$ in characteristic 00, then $\mathbb{U}(A_{E})=R_{F/\mathbf{Q}}\mathbb{G}_{m}$ and $E=kF$. If $A$ is ordinary in charactertistic $p$, then $E=k$, while if it is supersingular, then $E=kF_{0}$ if $k_{0}$ is the field of constants of $E$ and $F_{0}$ is the smallest (at most quadratic) extension of $k_{0}$ such that $\operatorname{End}^{0}(A_{F_{0}})$ is a quaternion algebra.
@@ 2. Technical lemmas
Since $CH^{1}_{(0)}(A)=L^{1}_{(0)}(A)$ generates $L^{*}_{(0)}(A)$ multiplicatively and since $CH^{1}_{(0)}(A)\to A^{1}_{\operatorname{num}}(A)$ is surjective, $\rho$ is surjective. On the other hand, O’Sullivan [26, Th. 6.1.1] has constructed a ring-theoretic section of the homomorphism $CH^{*}(A)\to A^{*}_{\operatorname{num}}(A)$ which sends $A^{1}_{\operatorname{num}}(A)$ into $CH^{1}_{(0)}(A)$, hence restricts to a ring-theoretic section $\sigma$ of $\rho$. But $CH^{1}_{(0)}(A)\to A^{1}_{\operatorname{num}}(A)$ is even an isomorphism, hence $\sigma$ is surjective. ∎
@@ 4.1. A crude category
When $f$ is a homomorphism, this follows from Theorem 1 since $[\gamma_{f}]=(\gamma_{f})_{*}([A])$. (See [23, 1.3] or [24, Th. 5.10] for a different proof.) In general, $f$ is the composition of a homomorphism and a translation, so it remains to handle the latter case. But if $a\in A(k)$, the translation defined by $a$ may be written as a composition
@@
As usual, we define $\operatorname{\mathbf{LMot}}_{\sim}^{\operatorname{eff}}(k)_{0}$ as the Karoubian envelope of $\operatorname{\mathbf{LCorr}}_{\sim}(k)_{0}$. As observed in [24, p. 672], the Chow-Künneth projectors of $A$ belong to $L^{*}(A\times A)$ as well as the Lefschetz isomorphisms of [22] and their inverses, which justifies the second claim already in $\operatorname{\mathbf{LMot}}_{\sim}^{\operatorname{eff}}(k)_{0}$. Note in particular that $\mathbb{L}\in\operatorname{\mathbf{LMot}}^{\operatorname{eff}}_{\sim}(k)_{0}$ by using an elliptic curve (which makes it unnecessary to involve projective spaces in its definition as in [23]), so that we can define $\operatorname{\mathbf{LMot}}_{\sim}(k)_{0}$ by $\otimes$-inverting $\mathbb{L}$. Its rigidity is checked on additive generators $Lh_{\sim}(A)$ in the usual way, using the diagonal of $A\times A$ to define unit and counit, and (4.2) is also checked as usual. ∎
@@ 4.4. Proof of Theorem 4
The existence of the Deninger-Murre Chow-Künneth projectors (Corollary 4.1) is the analogue of [25, I, Conj. (A)]. Reasoning as in [25, I, 2.5.4], the corresponding analogue to Murre’s filtration then verifies $F^{\nu}L^{j}(A)=\bigoplus_{s\geq\nu}L^{j}_{s}(A)$ for $0\leq\nu\leq j$ and $F^{j+1}L^{j}(A)=0$. Since $\operatorname{Ker}(L^{g}(A\times A)\to L^{g}_{\operatorname{num}}(A\times A))$ is a nilideal by [21], any other choice of Chow-Künneth projectors is conjugate to the canonical one under a self-correspondence $1+n$ with $n\sim_{\operatorname{num}}0$, as in [15, Lemma 5.4]. By (iii) and (iv) applied to $A\times A$, we have $n\in L^{g}_{(>0)}(A\times A)$, hence $nL^{j}_{(s)}(A)\subseteq\bigoplus_{t>s}L^{j}_{(t)}(A)$. This implies that the filtration $F^{\nu}L^{*}(A)$ does not depend on the choice of a Chow-Künneth decomposition, which is the analogue of Murre’s Conjecture (C). Finally, we get the analogue of Murre’s conjecture (B) (resp. (D)) from (iv) (resp. (iii)).
@@ 5.1. Centralisers in Tannakian categories
Recall [28, III.1] (see also [4, 5.3.1]) that one can define a $K$-linear Tannakian category $\mathcal{A}_{K}$ and a $Q$-linear $\otimes$-functor $(-)_{K}:\mathcal{A}\to\mathcal{A}_{K}$ such that $\mathcal{A}(A,B)\otimes_{Q}K\to\mathcal{A}_{K}(A_{K},B_{K})$ is an isomorphism for any $A,B\in\mathcal{A}$ and that $\omega$ extends canonically to a fibre functor on $\mathcal{A}_{K}$. This reduces us to the case where $Q=K$. The claim then follows from Tannakian theory, since $\mathcal{A}$ gets identified with $\operatorname{\mathbf{Rep}}_{K}(H)$. ∎
@@
$\rho^{2}(f)=\omega(u)^{-1}(1_{\omega(L)}\otimes\rho(f)^{\vee})\omega(u)\\ =\omega(u)^{-1}(1_{\omega(L)}\otimes(\omega(u)^{-1}(1_{\omega(L)}\otimes f^{\vee})\omega(u))^{\vee})\omega(u)\\ =\omega(u)^{-1}(1_{\omega(L)}\otimes(\omega(u)^{\vee}(1_{\omega(L^{\vee})}\otimes f)(\omega(u)^{-1})^{\vee}))\omega(u)\\ =\omega(u^{-1}u^{\vee})f\omega((u^{-1})^{\vee}u)$
@@
where the fibre product is with respect to the second projections (cf. [24, Def. 4.6]). Here, we use the isomorphisms (5.2) for $rM$ and $M\oplus N$ obtained by direct sums from those of $M$ and $N$.
@@ 5.2. Weight gradings
is surjective for all $r,n$ by [24, Prop. 3.4] (whose proof uses invariant theory). The claim now follows from Lemma 5.1.
@@ 5.3. An “easy” exactness criterion
(In other words, Condition H0 is sufficient in [17, Lemma C.1] when $G$ and $H$ are proreductive.)
@@ 5.4. Action of a profinite group
See also [19] for a different presentation.
@@
is exact at $G$ in general [^5] See however [19, Theorem 4.6].. This is true if $\mathcal{A}$ is semi-simple:
@@
In c), $\rho$ exists by [16, Prop. 5.3], which also gives the following recipe: for $S$ simple, $\rho(S)=S$ if $S\in\mathcal{A}^{0}$ and $\rho(S)=0$ otherwise. It suffices to check that this matches with the formula of the statement, which follows from b).
@@ 5.5. Proof of Theorem 5
Proposition 5.17 also applies to prove [10, Prop. 6.23 (a)] – except for the connectedness of $G(k_{s})$ which would require the “Hodge = absolute Hodge” conjecture –, and [2, 4.6, exemples]. It also applies to numerical Grothendieck motives under the standard conjecture D or, unconditionally, to this category restricted to numerical motives of abelian type in characteristic 00. Since the Tannakian group $G_{s}$ is not connected (see [24, Table 2 p. 655]), one sees that the expected connectedness of the motivic Galois group over separably closed fields cannot be hoped to be proven by purely formal arguments in the above style.
@@ 6. Remarks and questions
In particular, do the refined (Chow-)Künneth decompositions of [20, Th. 7.7.3]
@@
coincide in the two categories? This is true for $i=2$ by a generalisation of [20, Proof of Prop. 7.2.3]. However the answer is no in general, as pointed out by Peter O’Sullivan by the following simple argument: